OpenStudy (anonymous):

Solve: 3(x + 4) = 5(x − 2)

OpenStudy (anonymous):

3(x+4)=5(x-2) 3x+12=5x-10 5x-3x=12+10 2x=22 x=??

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

yes.. do you understand.>>>

OpenStudy (anonymous):

yes thnx